OpenStudy (aravindg):

5. A body of mass 2 kg is acted upon by two forces. A force of 8 Newton acts along positive direction of X-axis and another force of 6 Newton along positive direction of Y-axis. What is the acceleration of the body?

OpenStudy (anonymous):

simple 8-6=2

OpenStudy (aravindg):

how?

OpenStudy (anonymous):

bec forces are acting perpendicularly

OpenStudy (aravindg):

show step plzzz

OpenStudy (anonymous):

w8

OpenStudy (anonymous):

F=ma F(x)=8 -> a(x)=F(x)/m -> 8/2=4 F(y)=6 -> a(y)=F(y)/m -> 6/2=3 Net force is (a(x)^2+a(y)^2)^1/2 = (4*4+3*3)^1/2=25^1/2=5 Net acc. = 5 See - are you a blond?

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

thx